# What If Moderation Didn’t Mean Suppression? A Case for Personalized Content Transformation

### v2

@@ 3.3.1. A Spectrum of Triggers: From Phobias to Values
Beyond the relatively “easy” cases where restrictions are broadly accepted (schoenebeck2023online; klonick2017new), much of the tension over what should or should not be allowed in online spaces stems from the legitimate subjectivity of harm.
@@ 3.3.2. The Nuance of a Trigger: Granularity and Context-Dependence
This sensitivity sometimes extends to unrelated objects sharing similar visual patterns which then activate the same response. P_0_04 described experiencing her centipede phobia when seeing brain coral in his son’s ocean book because the coral’s pattern resembled centipede shapes.
Design Guideline 1 Enable users to define content sensitivities at their chosen level of specificity rather than constraining them to predetermined classification schemes.
@@ 3.4.1. When Platform Tools Fail and Lose Trust
Design Guideline 2 Build transparent systems where users understand when content is filtered and can easily override decisions, building trust through control.
@@ 3.4.2. The Demand for Personal Agency
Design Guideline 3 Implement intervention at the recipient level, rather than constraining the poster, ensuring one user’s safety needs do not restrict others’ expression.
@@ 3.5.1. The Blunt Instrument Problem
Design Guideline 4 Recognize that harmful and valuable content often coexist. Design systems that modify the experiential layer rather than removing content entirely.
@@ 3.6.1. Manual Labor of Safety
Design Guideline 5 Minimize the labor of safety by supporting adjustments that adapt to a user’s fluctuating sensitivities.
@@ 3.7. Context Determines Content Impact
Design Guideline 6 Design interventions that adapt to individual differences in how content modality and presentation affect distress.
@@ 4. The System: DIY-MOD
We designed DIY-MOD as an open source browser extension 555 The source code can be found at https://github.com/UMichHCI/diy_mod that creates a personalized content moderation layer between users and platform content. Consider Rex from our opening vignette. After installing DIY-MOD, he opens the extension popup. Through the conversational interface, he types: “I have an eating disorder that I’m trying to manage. Food pictures trigger cravings and obsessive thoughts I can’t control.” He then configures the filter with a high sensitivity level to apply only to images, setting the duration to ‘Always’.
Minutes later, Rex scrolls his Reddit’s popular feed: r/popular and encounters a post celebrating a community event, with an image showing people gathered around a table filled with food. DIY-MOD intercepts the posts (text and image) before it reaches the browser. The system’s backend analyzes the text and image against Rex’s filter and applies an appropriate transformation, such as using semantic inpainting to remove the food while preserving the social context of the gathering (Figure˜1b). A small “Modified by DIY-MOD” indicator appears (Figure˜1d). Rex can now safely engage with the post without being exposed to his triggers. The original post remains unchanged for everyone else; the transformation happens only in Rex’s browser.
This scenario illustrates DIY-MOD’s three core components. First, users create filters through natural language conversation or image upload using the extension popup. Second, platform adapter intercepts incoming content during natural browsing before it is handed off for rendering (piccardi2024reranking; kolluri2025alexandria). Third, our backend analyzes content using large vision-language models (LVLMs) and applies appropriate transformations and returns instructions and data to the adapter to render locally. We describe them in the following sections.
@@ 4.1. Filter Creation and Configuration
Users describe their sensitivities in natural language to DIY-MOD using their own words or by uploading example images. To capture the user’s specific needs, the system engages in conversational grounding (cho2020role; brennan2014grounding). The underlying language model is prompted to seek clarification when a description is ambiguous or underspecified, helping the user iteratively refine the filter. For instance, after a user uploads an image of jewelry, the system might ask: “Does this sensitivity apply to all jewelry or specific types?” Similarly, an initial filter for political content might prompt the system to ask: “Does this apply to all political topics, or specific ones like discussions of civilian casualties?”
@@
After establishing the filter description, users access three configuration controls (see Section˜C.1):
@@
The extension popup interface provides essential controls within limited screen space. Users manage their filters (e.g., custom time limits, additional metadata, or editing filter descriptions) by accessing the options page (options_page) of the extension. When creating filters, DIY-MOD can also detect potential overlaps and offer to modify existing filters.
@@ 4.2. Content Processing Client
The client employs platform-specific adapters that intercept network responses containing new content elements (piccardi2024reranking). This design keeps the filtering logic platform-independent, which we validated by implementing an adapter for Reddit. When adapters detect new posts, they extract text and image URLs before content is handed off for rendering. Only public content leaves the client and no user identifiers or profile information is transmitted to the server. The server analyzes this content against the user’s filters and returns personalized transformation instructions. The adapter then applies transformations to content within the post’s original layout before content becomes visible to the user. This client-side application of targeted modifications ensures transformations are applied at the recipient level (DG3) and preserves valuable surrounding elements within the content (DG4).
Transparency in Modified Content
DIY-MOD client marks all modified content with visible cues, displaying a “Modified by DIY-MOD” badge at the top right (Figure˜1d). When sensitive content dominates a post—making partial modification insufficient—the system displays personalized warning overlays that prevent unexpected exposure. Users can click through to view the original content if they so choose. The visible indicators and override options build trust by giving users control (DG2).
4.3. Server Architecture and Processing Pipeline
The DIY-MOD server handles both filter creation and content processing. For filter creation, it maintains stateful sessions with client chat interface. For content processing, it analyzes batches of posts from client adapters. When users browse supported platforms, the client adapter extracts content from API responses each containing a batch of posts (typically 25 for Reddit) and sends the content to the server. The server uses LLMs and LVLMs to evaluate the text and images in each post against the user’s active filters. When a match is found, the server initiates its transformation pipeline to apply an “intervention”. Each intervention represents a different way to modify sensitive content while preserving informational value. Section˜4.4 describes the range of interventions we explored, and Section˜4.5 details our two-stage pipeline for selecting the most appropriate one in real-time for each user and post.
The computational demands of this server processing pipeline are substantial. For example, processing a single Reddit batch can require 50 to 250 sequential external LVLM API calls that must complete within seconds to maintain usability. In our deployment, the critical path for processing an initial 25-post Reddit batch completed in roughly 5–15 seconds, with subsequent batches typically processed before users scrolled to them. We achieved real-time performance through asynchronous processing with multi-batch responses, content-based caching, and predictive prefetching. We have described these optimizations and additional design decisions in Appendix˜A.
4.4. The Intervention Palette
@@
Figure 2. Intervention palette demonstrating DIY-MOD’s three categories of transformation, applied to an original image (a). Semantic Modification[b, c] alters the content through techniques like (b) inpainting, which removes the trigger and (c) replacement with user-specified alternatives—trees, stars, cards in this case. Obfuscation[d] reduces trigger fidelity, shown here with blurring (d). Stylistic Alteration[e, f] changes the rendering style; a Studio Ghibli style animation is shown applied to the entire image (e) and selectively to only the trigger region (f). The transformations shown address the filter created by Rex (Section˜1), who is managing binge-eating disorder.
In contrast, semantic modifications prioritize perceptual smoothness. Inpainting(yu2018inpainting; li2024mat) as an example case, can seamlessly remove an object and reconstruct the background. This is a greater sacrifice of semantic fidelity than occlusion because it actively creates a new, plausible reality where the object never existed, rather than simply occluding part of the original. Another semantic intervention, visual euphemism, explores a different trade-off. It replaces a triggering object with a benign alternative, for instance, substituting spider with a leaf. While this also sacrifices semantic fidelity, the goal is to produce a visually coherent image that we hypothesize is less distressing to the user. To personalize this, users can add additional metadata to their filters or give examples to guide the visual euphemism choices, which may in turn alter the mood of the post.
Third, we also explore how altering an image’s rendering style could offer a middle ground. For example, reducing photorealism with artistic alteration, such as  Pointillism (clement1999neo) can lower the trigger fidelity of a sensitive region while preserving more structural information than occlusion. For these alterations, we borrow the aesthetic language of well-known artistic styles—Pointillism, Cubism, Impressionism, and Studio Ghibli style animation. We intended for each artistic style to not only reduce realism, but also act as an affective vessel that shapes how content is experienced. Impressionism, with its soft brushstrokes and diffused light, often evokes a sense of tranquility, distance, or dream-like calm. Studio Ghibli inspired style, with its rounded and whimsical rendering, lends even serious scenes a tone of cuteness and warmth. Cubism, which fragments forms in abstract planes, introduces a kind of conceptual distance, perhaps inviting interpretation over emotional reaction.
In DIY-MOD, we implemented a representative set of interventions that explore the broader design space of content transformation. Figure˜2 illustrates several examples from our image palette’s three main categories. We list all implemented interventions in Table˜1 & 2 . The range of transformation options presents an important question: which intervention is most appropriate for a specific user and post? Next, we detail the framework we developed to address this question.
@@ 4.5. Intervention Selection Framework
We developed a two-stage pipeline (Figure˜3) that evaluates interventions through the lens of individual user needs while maintaining real-time performance. Content that matches a user’s filters enters Stage 1 for pruning, and Stage 2 then generates and scores candidate transformations before selecting the best one.
@@ 4.5.1. Evaluation Criteria:
The three dimensions of semantic fidelity, trigger fidelity, and perceptual smoothness characterize the design space, but do not, on their own, determine which intervention is best suited for a given scenario. That choice requires translating these properties into outcomes relevant to a specific user’s safety and context. To guide this selection and evaluate outcomes, we operationalize a framework that decomposes intervention quality into four complementary dimensions, assessed through LLM-as-a-judge (zheng2023judging; liu2023gpteval):
@@
    Semantic Fidelity: Inherited directly from our design space, this assesses whether original facts and context are preserved without adding misinformation or hallucinations, a known risk in generative models (huang2025survey; gao2024aigcs).
@@
    Contextual Harm Risk: This dimension functions as a safeguard, assessing the degree to which a transformation might trivialize or distort themes that individuals or broader culture consider sensitive. For instance, applying Studio Ghibli style to war imagery or racial violence would would produce a high contextual-harm score. This risk score is subtracted from the overall intervention rating, so transformations that produce contextually inappropriate content are strongly penalized even if they score well on other dimensions.
@@ 4.5.2. Two-Stage Selection Pipeline
Stage 1: Predictive Pruning. The Pruner uses GPT-4o to perform multi-task inference(son2024multi) that first checks whether content matches any user filters. For matching content, it predicts which interventions would most likely succeed for that specific user. We provide the Pruner LVLM with detailed descriptions of each intervention including trade-offs, advantages, disadvantages, and ideal use cases. We also supply our four scoring dimensions and instruct the Pruner to evaluate how each intervention would score. This process approximates the full scoring process without generating actual transformations and produces a ranked list from which we select the top-$K$ candidates (typically $K=3$).
Stage 2: Generation and Scoring. DIY-MOD applies the top-$K$ candidates to the content, in parallel. Obfuscation techniques are applied locally using GroundingDino (liu2023grounding) and Python Pillow (clark2015pillow) library. For semantic modifications and stylistic alterations, we use Gemini’s asynchronous image generation model: gemini-2.0-flash-preview-image-generation (google_gemini2_0flash_2025). The Scorer uses GPT-4o-2024-08-06 to evaluate each candidate transformation against the evaluation dimensions (Section˜4.5.1). To mitigate potential biases inherent in LLM-as-a-judge systems where a model may favor its own outputs, we deliberately employ a different model for the scoring phase (GPT-4o) than for the generation phase (Gemini 2.0). We implement two guardrails in the scoring process. First, the Scorer receives both the original and transformed content, and is instructed to penalize misapplied interventions and failures. Second, we employ two-stage prompting (wang2024rl) that improve judge-VLM’s performance and results in more reliable and consistent evaluations. The Scorer first performs chain-of-thought analysis along our evaluation dimensions given the original content, transformed candidate, and user context. Based on this free-form response, it then produces a final numerical score. These scoring operations run in parallel across all $K$ candidates, and the highest-scoring intervention is selected.
@@
For text transformations and simple image obfuscations (e.g., Occlusion), the system sends lightweight instructions that are applied directly in the user’s browser using CSS. For other transformations, the generated content is served to the client. In all cases, we apply modification indicator badge (Section˜4.2) for transparency and store the resulting content in server cache (Section˜A.4).
@@ 5.1. Methods
Data Collection and Analysis: We transcribed all think-aloud data and interviews, then analyzed them using reflexive thematic analysis (braun2019reflecting) to identify patterns in user experiences and system interactions. During consent, we explicitly warned participants they might encounter content matching their stated sensitivities and emphasized they could pause, skip content, or withdraw at any point.
@@ 6.1. Method
As participants scrolled through the dual feed, they encountered two distinct, modified versions of each post with text and image transformation applied ( Figure˜4). For each pair, they were asked to select the version that better met their safety needs and explain their reasoning. The two modifications were chosen systematically: one was the intervention ranked highest by DIY-MOD’s selection framework, while the other was randomly selected from the framework’s other high-ranking alternatives (ranks 2-5). To mitigate positional bias, the on-screen placement (left or right) of these two versions was randomized for each post. While the primary task involved comparing the two modified versions, the original unmodified content was also available for context. Participants could request to view it or have it described verbally if needed. The study lasted approximately 30 minutes for each participant.

### v3

@@ 3.3.1. A Spectrum of Triggers: From Phobias to Values
Beyond the relatively “easy” cases where restrictions are broadly accepted (Schoenebeck et al., 2023; Klonick, 2017), much of the tension over what should or should not be allowed in online spaces stems from the legitimate subjectivity of harm.
@@ 3.3.2. The Nuance of a Trigger: Granularity and Context-Dependence
This sensitivity sometimes extends to unrelated objects sharing similar visual patterns which then activate the same response. P_0_04 described experiencing her centipede phobia when seeing brain coral in her son’s ocean book because the coral’s pattern resembled centipede shapes.
Design Guideline 1
Enable users to define content sensitivities at their chosen level of specificity rather than constraining them to predetermined classification schemes.
@@ 3.4.1. When Platform Tools Fail and Lose Trust
Design Guideline 2
Build transparent systems where users understand when content is filtered and can easily override decisions, building trust through control.
@@ 3.4.2. The Demand for Personal Agency
Design Guideline 3
Implement intervention at the recipient level, rather than constraining the poster, ensuring one user’s safety needs do not restrict others’ expression.
@@ 3.5.1. The Blunt Instrument Problem
Design Guideline 4
Recognize that harmful and valuable content often coexist. Design systems that modify the experiential layer rather than removing content entirely.
@@ 3.6.1. Manual Labor of Safety
Design Guideline 5
Minimize the labor of safety by supporting adjustments that adapt to a user’s fluctuating sensitivities.
@@ 3.7. Context Determines Content Impact
Design Guideline 6
Design interventions that adapt to individual differences in how content modality and presentation affect distress.
@@ 4. The System: DIY-MOD
We designed DIY-MOD as an open source browser extension 555 The source code can be found at https://github.com/UMichHCI/diymod that creates a personalized content moderation layer between users and platform content. Consider Rex from our opening vignette. After installing DIY-MOD, he opens the extension popup. Through the conversational interface, he types: “I have an eating disorder that I’m trying to manage. Food pictures trigger cravings and obsessive thoughts I can’t control.” He then configures the filter with a high sensitivity level to apply only to images, setting the duration to ‘Always’.
Minutes later, Rex scrolls his Reddit’s popular feed: r/popular and encounters a post celebrating a community event, with an image showing people gathered around a table filled with food. DIY-MOD intercepts the posts (text and image) before it reaches the browser. The system’s backend analyzes the text and image against Rex’s filter and applies an appropriate transformation, such as using semantic inpainting to remove the food while preserving the social context of the gathering (Figure 1b). A small “Modified by DIY-MOD” indicator appears (Figure 1d). Rex can now safely engage with the post without being exposed to his triggers. The original post remains unchanged for everyone else; the transformation happens only in Rex’s browser.
This scenario illustrates DIY-MOD’s three core components. First, users create filters through natural language conversation or image upload using the extension popup. Second, platform adapter intercepts incoming content during natural browsing before it is handed off for rendering (Piccardi et al., 2024; Kolluri et al., 2025). Third, our backend analyzes content using large vision-language models (LVLMs) and applies appropriate transformations and returns instructions and data to the adapter to render locally. We describe them in the following sections.
@@ 4.1. Filter Creation and Configuration
Users describe their sensitivities in natural language to DIY-MOD using their own words or by uploading example images. To capture the user’s specific needs, the system engages in conversational grounding (Cho and Rader, 2020; Brennan, 2014). The underlying language model is prompted to seek clarification when a description is ambiguous or underspecified, helping the user iteratively refine the filter. For instance, after a user uploads an image of jewelry, the system might ask: “Does this sensitivity apply to all jewelry or specific types?” Similarly, an initial filter for political content might prompt the system to ask: “Does this apply to all political topics, or specific ones like discussions of civilian casualties?”
@@
After establishing the filter description, users access three configuration controls (see Appendix B):
@@
The extension popup interface provides essential controls within limited screen space. Users manage their filters (e.g., custom time limits, additional metadata, or editing filter descriptions) by accessing the options page (for Developers Team, 2012) of the extension. When creating filters, DIY-MOD can also detect potential overlaps and offer to modify existing filters.
@@ 4.2. Content Processing Client
The client employs platform-specific adapters that intercept network responses containing new content elements (Piccardi et al., 2024). This design keeps the filtering logic platform-independent, which we validated by implementing an adapter for Reddit. When adapters detect new posts, they extract text and image URLs before content is handed off for rendering. Only public content leaves the client and no user identifiers or profile information is transmitted to the server. The server analyzes this content against the user’s filters and returns personalized transformation instructions. The adapter then applies transformations to content within the post’s original layout before content becomes visible to the user. This client-side application of targeted modifications ensures transformations are applied at the recipient level (DG3) and preserves valuable surrounding elements within the content (DG4).
@@
Figure 2. Intervention palette demonstrating DIY-MOD’s three categories of transformation, applied to an original image (a). Semantic Modification[b, c] alters the content through techniques like (b) inpainting, which removes the trigger and (c) replacement with user-specified alternatives—trees, stars, cards in this case. Obfuscation[d] reduces trigger fidelity, shown here with blurring (d). Stylistic Alteration[e, f] changes the rendering style; a Studio Ghibli style animation is shown applied to the entire image (e) and selectively to only the trigger region (f). The transformations shown address the filter created by Rex (Section 1), who is managing binge-eating disorder.
Transparency in Modified Content
DIY-MOD client marks all modified content with visible cues, displaying a “Modified by DIY-MOD” badge at the top right (Figure 1d). When sensitive content dominates a post—making partial modification insufficient—the system displays personalized warning overlays that prevent unexpected exposure. Users can click through to view the original content if they so choose. The visible indicators and override options build trust by giving users control (DG2).
4.3. Server Architecture and Processing Pipeline
The DIY-MOD server handles both filter creation and content processing. For filter creation, it maintains stateful sessions with client chat interface. For content processing, it analyzes batches of posts from client adapters. When users browse supported platforms, the client adapter extracts content from API responses each containing a batch of posts (typically 25 for Reddit) and sends the content to the server. The server uses LLMs and LVLMs to evaluate the text and images in each post against the user’s active filters. When a match is found, the server initiates its transformation pipeline to apply an “intervention”. Each intervention represents a different way to modify sensitive content while preserving informational value. Section 4.4 describes the range of interventions we explored, and Section 4.5 details our two-stage pipeline for selecting the most appropriate one in real-time for each user and post.
The computational demands of this server processing pipeline are substantial. For example, processing a single Reddit batch can require 50 to 250 sequential external LVLM API calls that must complete within seconds to maintain usability. In our deployment, the critical path for processing an initial 25-post Reddit batch completed in roughly 5–15 seconds, with subsequent batches typically processed before users scrolled to them. We achieved real-time performance through asynchronous processing with multi-batch responses, content-based caching, and predictive prefetching. We have described these optimizations and additional design decisions in Appendix A.
4.4. The Intervention Palette
@@
In contrast, semantic modifications prioritize perceptual smoothness. Inpainting(Yu et al., 2018; Li et al., 2022) as an example case, can seamlessly remove an object and reconstruct the background. This is a greater sacrifice of semantic fidelity than occlusion because it actively creates a new, plausible reality where the object never existed, rather than simply occluding part of the original. Another semantic intervention, visual euphemism, explores a different trade-off. It replaces a triggering object with a benign alternative, for instance, substituting spider with a leaf. While this also sacrifices semantic fidelity, the goal is to produce a visually coherent image that we hypothesize is less distressing to the user. To personalize this, users can add additional metadata to their filters or give examples to guide the visual euphemism choices, which may in turn alter the mood of the post.
Third, we also explore how altering an image’s rendering style could offer a middle ground. For example, reducing photorealism with artistic alteration, such as  Pointillism (Clement and Houzé, 1999) can lower the trigger fidelity of a sensitive region while preserving more structural information than occlusion. For these alterations, we borrow the aesthetic language of well-known artistic styles—Pointillism, Cubism, Impressionism, and Studio Ghibli style animation. We intended for each artistic style to not only reduce realism, but also act as an affective vessel that shapes how content is experienced. Impressionism, with its soft brushstrokes and diffused light, often evokes a sense of tranquility, distance, or dream-like calm. Studio Ghibli inspired style, with its rounded and whimsical rendering, lends even serious scenes a tone of cuteness and warmth. Cubism, which fragments forms in abstract planes, introduces a kind of conceptual distance, perhaps inviting interpretation over emotional reaction.
@@
In DIY-MOD, we implemented a representative set of interventions that explore the broader design space of content transformation. Figure 2 illustrates several examples from our image palette’s three main categories. We list all implemented interventions in Table 1 and 2 . The range of transformation options presents an important question: which intervention is most appropriate for a specific user and post? Next, we detail the framework we developed to address this question.
@@ 4.5. Intervention Selection Framework
We developed a two-stage pipeline (Figure 3) that evaluates interventions through the lens of individual user needs while maintaining real-time performance. Content that matches a user’s filters enters Stage 1 for pruning, and Stage 2 then generates and scores candidate transformations before selecting the best one.
@@ 4.5.1. Evaluation Criteria:
The three dimensions of semantic fidelity, trigger fidelity, and perceptual smoothness characterize the design space, but do not, on their own, determine which intervention is best suited for a given scenario. That choice requires translating these properties into outcomes relevant to a specific user’s safety and context. To guide this selection and evaluate outcomes, we operationalize a framework that decomposes intervention quality into four complementary dimensions, assessed through LLM-as-a-judge (Zheng et al., 2023; Liu et al., 2023b):
@@
    Semantic Fidelity: Inherited directly from our design space, this assesses whether original facts and context are preserved without adding misinformation or hallucinations, a known risk in generative models (Huang et al., 2025; Gao et al., 2024).
@@
    Contextual Harm Risk: This dimension functions as a safeguard, assessing the degree to which a transformation might trivialize or distort themes that individuals or broader culture consider sensitive. For instance, applying Studio Ghibli style to war imagery or racial violence would produce a high contextual-harm score. This risk score is subtracted from the overall intervention rating, so transformations that produce contextually inappropriate content are strongly penalized even if they score well on other dimensions.
@@ 4.5.2. Two-Stage Selection Pipeline
Stage 1: Predictive Pruning. The Pruner uses GPT-4o to perform multi-task inference(Son et al., 2024) that first checks whether content matches any user filters. For matching content, it predicts which interventions would most likely succeed for that specific user. We provide the Pruner LVLM with detailed descriptions of each intervention including trade-offs, advantages, disadvantages, and ideal use cases. We also supply our four scoring dimensions and instruct the Pruner to evaluate how each intervention would score. This process approximates the full scoring process without generating actual transformations and produces a ranked list from which we select the top-$K$ candidates (typically $K=3$).
Stage 2: Generation and Scoring. DIY-MOD applies the top-$K$ candidates to the content, in parallel. Obfuscation techniques are applied locally using GroundingDino (Liu et al., 2023a) and Python Pillow (Clark, 2015) library. For semantic modifications and stylistic alterations, we use Gemini’s asynchronous image generation model: gemini-2.0-flash-preview-image-generation (Google, 2025). The Scorer uses GPT-4o-2024-08-06 to evaluate each candidate transformation against the evaluation dimensions (Section 4.5.1). To mitigate potential biases inherent in LLM-as-a-judge systems where a model may favor its own outputs, we deliberately employ a different model for the scoring phase (GPT-4o) than for the generation phase (Gemini 2.0). We implement two guardrails in the scoring process. First, the Scorer receives both the original and transformed content, and is instructed to penalize misapplied interventions and failures. Second, we employ two-stage prompting (Wang et al., 2024) that improve judge-VLM’s performance and results in more reliable and consistent evaluations. The Scorer first performs chain-of-thought analysis along our evaluation dimensions given the original content, transformed candidate, and user context. Based on this free-form response, it then produces a final numerical score. These scoring operations run in parallel across all $K$ candidates, and the highest-scoring intervention is selected.
@@
For text transformations and simple image obfuscations (e.g., Occlusion), the system sends lightweight instructions that are applied directly in the user’s browser using CSS. For other transformations, the generated content is served to the client. In all cases, we apply modification indicator badge (Section 4.2) for transparency and store the resulting content in server cache (Section A.4).
@@ 5.1. Methods
Data Collection and Analysis: We transcribed all think-aloud data and interviews, then analyzed them using reflexive thematic analysis (Braun and Clarke, 2019) to identify patterns in user experiences and system interactions. During consent, we explicitly warned participants they might encounter content matching their stated sensitivities and emphasized they could pause, skip content, or withdraw at any point.
@@ 6.1. Method
As participants scrolled through the dual feed, they encountered two distinct, modified versions of each post with text and image transformation applied (Figure 4). For each pair, they were asked to select the version that better met their safety needs and explain their reasoning. The two modifications were chosen systematically: one was the intervention ranked highest by DIY-MOD’s selection framework, while the other was randomly selected from the framework’s other high-ranking alternatives (ranks 2-5). To mitigate positional bias, the on-screen placement (left or right) of these two versions was randomized for each post. While the primary task involved comparing the two modified versions, the original unmodified content was also available for context. Participants could request to view it or have it described verbally if needed. The study lasted approximately 30 minutes for each participant.